Suppose that a firm would like to maintain the current debt-to-equity after it's growth. What part of the external funds needed will it raise in equity?

what does this mean?
Answer #1

The external funds shall be raised if the retained earnings could not provide enough cash flow for expansion or other investment. This means that while maintaining the capital structure, if there is a need for external financing, then only external funds shall be raised.

An example would provide a better explaination:
For example a company has a debt to equity ratio of 1:2 and funds needed are $900

The retained earnings are $300.

Therefore the debt shall be 1/3*$900=$300

Equity required=2/3*$900=$600

External equity needed=$600-$300=$300